OpenStudy (anonymous):

WHAT IS THE RANGE OF F(X)=25X^2 -30X+16

OpenStudy (anonymous):

range from the second coordinate of the vertex up to infinity first coordinate of the vertex is \(-\frac{b}{2a}\) second coordinate of the vertex is what you get when you replace \(x\) by the first coordinate

OpenStudy (anonymous):

in your case the first coordinate of the vertex is \[-\frac{-30}{2\times 25}=\frac{30}{50}=0.6\] so your job is to replace \(x\) by \(0.6\) so see what you get